# GPTKB: Comprehensively Materializing Factual LLM Knowledge

### v3

@@ 6.1 Lessons for LLM epistemiology
The notion and definition of LLM knowledge itself is controversial Fierro et al. (2024). We adopt here the term knowledge base because it is common in the field, but would consider the term belief base equally appropriate. In terms of the definitions of knowledge that are introduced by Fierro et al. (2024), we find that our output falls into the minimalistic sui-generis (g-knowledge) category, as our KB contains false as well as inconsistent triples. Beyond the reported errors, e.g., in many cases, spouse relations are not symmetrically represented.
@@
We observe a high complementarity to existing resources, with 63% of generated entities not being present in Wikidata. Particularly, from Figure 3, GPTKB’s entities remain web-verifiable even in deeper layers, about 70% of the entities are web-verifiable in layer 10, even though the percentage of entities found in Wikidata drops to a little over 20%. Apart from named-entities, some particularly noteworthy complementary slices of LLM knowledge concern digital media artifacts, art periods, and people’s hobbies.
@@
Our materialized resource enables a range of analyses about factual LLM knowledge, which for space reasons we only skim here: 1) Accuracy: See Sec. 5.1 and paragraph above. 2) LLM bias: See Sec. 5.2. 3) Timeliness: Exemplarily, we collect the most frequently mentioned years, and observe that there is sharp drop after 2023, which matches the knowledge cutoff of the LLM (see plot in Appendix D). 4) Subject-wise consistency: Our KB still contains many duplicates, e.g., The Elbe River, River Elbe, Elbe River, river Elbe, Elbe. Studying their triples gives insights into consistency, e.g., we observe a significant semantic overlap, but also frequent different wordings (e.g., wildlife - various fish species / fish species), and minor factual deviations (e.g., length 3x 1094km, 2x 1091km). 5) Structural consistency: An interesting observation concerns the difficulty that LLMs have with inverse relations Berglund et al. (2024). We can observe this too, for instance, out of 318K spouse triples, only 8K are symmetric, and out of 61K parentCompany triples, only 6K are mirrored in subsidiary triples.
6.2 Lessons for KB construction
The GPTKB prototype reveals several important lessons for LLM-based KB construction. In particular, we find it notable that building such a large KB was possible so quickly, with a relatively small model.
@@ 6.2 Lessons for KB construction
The GPTKB approach deviates from traditional Wikimedia- and data-integration focused KB construction, and although its precision still needs improvement, it stands out with its potential for cost efficiency. In a back-of-the-envelope calculation, Paulheim (2018) estimated the cost per triple for existing manual KB construction projects at \$2-6, and for existing automated KB construction projects at \$0.01-0.15. In contrast, in our prototypical execution, the cost is just \$0.0001 per correct triple (\$3.5K/33M), i.e., $>$100x less than with previous methods.
@@ 6.3 Other LLMs
We performed a parallel exploration using Llama-3.1-70B-Instruct on local HPC hardware, and while accuracy is higher (69% of triples verified true in a test run), this did not scale. We also envisioned a run using the strongest publicly available LLM, GPT-4o (80% triples verified true), however, by its 15x higher API cost, and its estimated 25x larger parameter set, assuming knowledge is roughly proportional to parameters, at about \$825K, this is beyond our budget. See Appendix E.
@@ 7 Conclusion
Our work is a landmark for two fields: For NLP, for the first time, it provides comprehensive insights into what LLMs know (or believe). For the Semantic Web, it shows novel ways forward for the long stale topic of open-domain KB construction. GPTKB is accessible at https://gptkb.org.
@@ 8 Limitations
As already observed by Petroni et al. (2019), prompt formulations heavily influence LLM responses, and hence may bias the resulting KB. GPTKB presents one way of materializing LLM knowledge into structured format, but different prompts could give different output. It is also important to observe that LLMs operate inherently probabilistically, while our output (and most web-scale KBs) are binary. Any process translating from probabilistic to binary implicitly encodes a choice between precision and recall, and we do so by the triple counts given as guidance in the elicitation prompt. In our view, such a biasing is hard to avoid, and also not solved by instruction-free sample-only (few-shot) prompts as proposed by Wu et al. (2024), as also examples bias the responses, both in terms of quantity, topics, and vocabulary.
Our exemplary run presents a proof of concept for materializing LLM knowledge. While this knowledge has slices complementary to existing KB knowledge, so far our experiments on verifiability does not satisfy the precision requirements of most downstream use cases, nor is its precision even accurately established, because it contains many long-tail triples without web sources that enable confirming or refuting their truth. The knowledge in our KB can intrinsically not be sourced back to any specific document and contributor. Also, as we do not prescribe any standardized vocabulary, and only apply lightweight consolidation, our KB does not yet possess the consistency of existing projects. If you consider using a KB in production, for the time being, consider an alternative like Wikidata.
Our materialized KB is based on a closed-source LLM that also does not come with guarantees regarding persistent servicing. In the past, similar models have been discontinued, hence, long-term reproducibility is not guaranteed. An execution based on an open model is planned.
@@ Acknowledgment
We thank Markus Krötzsch and Gerhard Weikum for comments on this work.
@@ Appendix A Further KB Content
Figures 5 and 6 provide further examples of GPTKB content. In the first case, all triples appear correct, but the predicate famousFor dominates the entity. In second case, most triples are correct, but the spouse is made up. In the online OpenReview appendix to this submission, we provide a larger sample of 10K triples.
@@
Appendix B Class distribution by BFS-level
Figure 7 shows the most frequent classes by BFS-level. As one can see, diversity increases towards lower BFS-levels, and some classes, like awards, only emerge later.
Appendix C Prompts
@@
Appendix D Analysis of temporal cutoff
@@
Appendix E Comparison of LLMs
GPT Llama GPT
4o-mini* 3.1 70B 4o
Web-verified triples 0.38 0.69 0.78
Entities on Wikidata 0.78 0.83 0.88
Web-verified entities 0.80 0.95 0.98
Appendix F Sources for KB comparison
Sources of KB sizes in the comparison presented in Table 2: Wikidata: See https://www.wikidata.org/wiki/Wikidata:Statistics and https://grafana.wikimedia.org/d/000000175/wikidata-datamodel-statements. Wikidata5m: Wang et al. (2021). Yago4.5: Suchanek et al. (2024). DBpedia: English version as per Table 2 in Lehmann et al. (2015). NELL: As per Figure 5 (left) in Mitchell et al. (2018). Reverb: As per Lin et al. (2012) and https://web.archive.org/web/20220307185343/https://openie.allenai.org/.

### v4

@@ 6.1 Lessons for LLM epistemology
The notion and definition of LLM knowledge itself is controversial Fierro et al. (2024). We adopt here the term knowledge base because it is common in the field, but would consider the term belief base equally appropriate. In terms of the definitions of knowledge that are introduced by Fierro et al. (2024), we find that our output falls into the minimalistic sui-generis (g-knowledge) category, as our KB contains false as well as inconsistent statements. Beyond the reported errors, e.g., in many cases, spouse relations are not symmetrically represented (see “enabled analyses” below).
@@
Knowledge consistency
To assess the consistency of LLM knowledge, we repeated the elicitation 100 times on the entity Vannevar Bush. From the perspective of knowledge volume consistency, the results primarily fall into two clusters: one comprising 52 runs with a mean of 21 triples (std = 1), and another with 32 runs averaging 38 triples (std = 6). The remaining 10 runs failed to parse. In terms of knowledge content consistency, within the first cluster, we observe 1,116 total triples and 79 unique ones, i.e., each triple is shared on average by 14 runs. The average overlap between the sets measured using an exact match set intersection is 0.67. Semantically, these numbers represent a lower bound only, as responses sometimes contain paraphrases.
@@
We observe a high complementarity to existing resources, with 63% of generated entities not being present in Wikidata. Particularly, from Figure 3, GPTKB’s entities remain web-verifiable even in deeper layers, about 70% of the entities are web-verifiable in layer 10, even though the percentage of entities found in Wikidata drops to a little over 20%. Some particularly noteworthy complementary slices of LLM knowledge concern digital media artifacts, art periods, and people’s hobbies.
@@
Our materialized resource enables a range of analyses about factual LLM knowledge, which for space reasons we only skim here: 1) Accuracy: See Sec. 5.1 and paragraph above. 2) LLM bias: See Sec. 5.2. 3) Timeliness: Exemplarily, we collect the most frequently mentioned years, and observe that there is sharp drop after 2023, which matches the knowledge cutoff of the LLM (see plot in Appendix C). 4) Subject-wise consistency: Our KB still contains many duplicates, e.g., The Elbe River, River Elbe, Elbe River, river Elbe, Elbe. Studying their triples gives insights into consistency, e.g., we observe a significant semantic overlap, but also frequent different wordings (e.g., wildlife - various fish species / fish species), and minor factual deviations (e.g., length 3x 1094km, 2x 1091km). 5) Structural consistency: An interesting observation concerns the difficulty that LLMs have with inverse relations Berglund et al. (2024). We can observe this too, for instance, out of 318K spouse triples, only 8K are symmetric, and out of 61K parent_company triples, only 6K are mirrored in subsidiary triples.
6.2 Lessons for knowledge materialization
The GPTKB prototype reveals several important lessons for knowledge materialization. In particular, we find it notable that building such a large KB was possible so quickly, with a relatively small model.
@@ 6.2 Lessons for knowledge materialization
The GPTKB approach deviates from traditional Wikimedia- and data-integration focused KB construction, and although its precision still needs improvement, it stands out with its potential for cost efficiency. In a back-of-the-envelope calculation, Paulheim (2018) estimated the cost per triple for existing manual KB construction projects at \$2-6, and for existing automated KB construction projects at \$0.01-0.15. In contrast, in our prototypical execution, the API cost is just \$0.0001 per correct triple (\$3.5K/33M), i.e., $>$100x less than with previous methods.
@@ 6.3 Other LLMs
We performed a parallel exploration using Llama-3.1-70B-Instruct on local HPC hardware, and while accuracy is higher (69% of triples verified true in a test run), this did not scale. We also envisioned a run using the strongest publicly available LLM, GPT-4o (80% triples verified true), however, by its 15x higher API cost, and its estimated 25x larger parameter set, assuming knowledge is roughly proportional to parameters, at about \$825K, this is beyond our budget.
GPT Llama GPT
4o-mini* 3.1 70B 4o
Web-verified triples 0.38 0.69 0.78
Entities on Wikidata 0.78 0.83 0.88
Web-verified entities 0.80 0.95 0.98
@@ 7 Conclusion
We propose a novel methodology to recursively materialize LLM knowledge. For the first time in NLP and LLM research, our work provides comprehensive insights into what LLMs know (or believe). Our resource is accessible at https://gptkb.org.
@@ 8 Limitations
Prompt dependency
As already observed by Petroni et al. (2019), prompt formulations influence LLM responses, and hence the resulting KB. GPTKB presents one way of materializing LLM knowledge into structured format, but better prompts may exist. In line with Petroni et al. (2019), methods like ours therefore always inform about lower bounds (on LLM factual knowledge, in our case).
Reproducibility
Our materialized KB is based on a closed-source LLM that does not come with guarantees regarding persistent servicing, and in the past, similar services have been discontinued, hence, long-term reproducibility is not guaranteed. At the same time, our resource therefore has especial archival value. An execution based on an open model is planned.
@@ Acknowledgment
We thank Markus Krötzsch and Gerhard Weikum for comments on this work, and Moritz Müller for help with the website.
@@ Appendix A Further KB Content
Figures 7 and 8 provide further examples of GPTKB content. In the first case, all triples appear correct, but the predicate famous_for dominates the entity. In second case, most triples are correct, but the spouse is made up. In the online OpenReview appendix to this submission, we provide a larger sample of 10K triples.
@@
Appendix B Prompts
@@
Appendix C Analysis of temporal cutoff
@@
Appendix D Comparison of GPTKB with other KBs
Table 2 gives a comparison of GPTKB to major KB projects. Sources: Wikidata: See https://www.wikidata.org/wiki/Wikidata:Statistics and https://grafana.wikimedia.org/d/000000175/wikidata-datamodel-statements. Wikidata5m: Wang et al. (2021). Yago4.5: Suchanek et al. (2024). DBpedia: English version as per Table 2 in Lehmann et al. (2015). NELL: As per Figure 5 (left) in Mitchell et al. (2018). Reverb: As per Lin et al. (2012) and https://web.archive.org/web/20220307185343/https://openie.allenai.org/.